OpenStudy (anonymous):

find the sum of the convergent series 3-1+1/3-1/9+...

OpenStudy (stamp):

formula

OpenStudy (raden):

S = a/(1-r) with a=3, r=-1/3

OpenStudy (anonymous):

Thank you very much. That helped a lot.

OpenStudy (raden):

you're welcome